When the sum of a number and 10 is doubled the result is the same as 8 minus twice the number. find the unknown number.
Answer by Mathtut(3670)

let x be the number
:
2(x+10)=8-2x
:
2x+20=8-2x
:
4x=-12
:
x=-3
: